# MODEM: A Morton-Order Degradation Estimation Mechanism for Adverse Weather Image Recovery

### v1

@@ 1 Introduction
Computer vision systems are increasingly integral to critical applications, such as autonomous driving [41, 1] and intelligent surveillance [39, 59], demanding reliable performance in diverse environments. However, their effectiveness deteriorates significantly under adverse weather conditions, such as rain [11, 32, 46, 62, 65], haze [63, 64, 16, 51, 60, 70, 71], and snow [28, 35, 48, 73], which introduce complex visual artifacts and obscure critical scene information. Thus, effectively restoring clear images from such weather-degraded inputs is essential to boost the robustness and safety of modern computer vision technologies in real-world deployments.
Early task-specific methods [2, 4, 18, 49, 50, 22, 24] largely rely on physical models or hand-crafted statistical priors tailored to individual weather phenomena. Due to the limited feature representation capabilities, these schemes are brittle in the face of complex scenes or deviations from assumed degradation patterns. With the advent of deep learning, numerous deep networks have achieved impressive performance by training on large-scale datasets for specific tasks such as image deraining [11, 32, 46, 62, 65, 5, 31, 58, 70], dehazing [63, 64, 16, 51, 60, 70, 71], or desnowing [28, 35, 48, 73]. These models excel at implicitly learning the inverse mapping through extensive supervision. But their highly specialized nature necessitates separate models for each weather condition, severely limiting scalability and practical deployment in unconstrained environments.
@@
Motivated by the above insight, we propose a novel Morton-Order Degradation Estimation Mechanism (MODEM) that estimates degradation state via SSM and guides adaptive restoration in degraded images. At the heart of MODEM lies the Morton-Order 2D-Selective-Scan Module (MOS2D), which integrates SSM with Morton-coded spatial traversal. Unlike conventional raster scanning, the Morton-order scan follows a hierarchical, locality-preserving sequence [40, 25, 43, 23] that enables structured and efficient modeling of both local and long-range dependencies, as shown in Fig. 1(d). To complement MOS2D, we further introduce a Dual Degradation Estimation Module (DDEM), designed to extract two complementary forms of degradation information from the input: (i) a global degradation representation that encapsulates high-level weather characteristics such as type and severity, and (ii) a set of spatially adaptive kernels that encode local degradation structures and variations. These dual degradation representations are then utilized to dynamically modulate the restoration process. Specifically, the global representation adaptively influences feature transformations within MODEM, while the spatially adaptive kernels guide the matrix construction within MOS2D, refining the spatial dependencies captured along the Morton-order sequence. This dual-modulation strategy equips the restoration network with both global awareness and local sensitivity, thereby delivering more precise and effective restoration. Our primary contributions can be summarized as follows:
@@ 2 Related Work
Raindrop Removal. Early raindrop removal efforts addressed specific scenarios [10] or adherent raindrops in videos [67]. Deep learning subsequently provided more generalizable solutions, including learning from synthetic photorealistic data [17], using attention-guided GANs for realistic inpainting [46], and dedicated networks for visibility through raindrop-covered windows [47]. General restoration architectures like Dual Residual Networks [33] are also applicable, with advanced dual attention-in-attention models [72] tackling joint rain streak and raindrop removal.
@@
Snow Removal. While specific traditional priors for snow removal are less common compared to other weather conditions, principles from general image restoration were often adapted. Deep learning solutions for desnowing have gained traction, often involving networks sensitive to image context [35] or explicitly leveraging semantic segmentation and depth estimation as priors [73]. Techniques inspired by classical methods like matrix decomposition have also been integrated into deep frameworks [48], alongside methods developed for online processing of surveillance videos [28].
@@ 3 Bridging Image Degradation Estimation and State Space Modeling
By comparing Eq. 1 with Eq. 3, we interpret SSM components as degradation estimators, as illustrated in Fig. 2. The term $CAh_{k-1}$ conceptually models the long-range, spatially propagated degradation context, as shown in Fig. 3(b). The previous hidden state $h_{k-1}$ summarizes accumulated degradation cues (e.g., overall haze level, rain intensity) along the Morton-order scan path. The state transition matrix $A$ then models how these summarized characteristics evolve or persist as the scan progresses spatially (e.g., the smooth attenuation of visibility due to widespread haze or the consistent statistical properties of rain streaks over a larger area.). It essentially dictates “how the long-range state evolves over pixels” as noted in Fig. 2. The output matrix $C$ then translates this evolved context $Ah_{k-1}$ into a contribution to the current output $y_{k}$, reflecting broader, non-local degradation patterns.
Concurrently, the $CBx_{k}$ term accounts for the impact of the immediate, local input features $x_{k}$, as visualized in Fig. 3(c). The input matrix $B$ determines “how the local information affects the current state” (Fig. 2), allowing features from $x_{k}$ (e.g., local degradation artifacts like dense fog patches, or crucial local image content like fine textures/sharp edges) to directly influence the current hidden state $h_{k}$. This enables the SSM to react to fine-grained local evidence, with $C$ projecting this locally-informed component into $y_{k}$ for specific adjustments based on immediate observations.
Figure 4: (a) Overall architecture of MODEM. (b) The DDEM for extracting global descriptor $Z_{0}$ and adaptive degradation kernel $Z_{1}$ degradation priors. (c) The MDSL incorporating the core MOS2D module (d) within a residual block. The blue-colored components indicate elements exclusive to the first training stage. $N$, $M_{1}$, $M_{2}$ denote the number of the corresponding module, respectively.
@@ 4 Morton-Order Degradation Estimation Mechanism
The Morton-Order Degradation Estimation Mechanism (MODEM), depicted in Fig. 4(a), employs a two-stage training strategy. During the first stage, the Dual Degradation Estimation Module (DDEM), shown in Fig. 4(b), utilizes both the degraded image $I_{\text{LQ}}$ and its corresponding ground-truth $I_{\text{GT}}$ to learn an accurate mapping to degradation priors: a global descriptor $Z_{0}$ and adaptive degradation kernel $Z_{1}$. In the second stage, DDEM processes only the degraded input $I_{\text{LQ}}$, with its estimation of $Z_{0}$ and $Z_{1}$ being supervised by the frozen first stage. The priors, $Z_{0}$ and $Z_{1}$, are injected into MODEM’s main restoration backbone in Fig. 4(a). This backbone is comprised of $N$ stacked MOS2D Degradation-Aware Layers (MDSLs), detailed in Fig. 4(c). Each MDSL leverages the Morton-Order 2D-Selective-Scan module (MOS2D), shown in Fig. 4(d) and Fig. 5, to adaptively modulate features based on the degradation priors $Z_{0}$ and $Z_{1}$.
@@ 4.1 Dual Degradation Estimation Module (DDEM)
The Dual Degradation Estimation Module (DDEM), shown in Fig. 4(b), extracts global degradation descriptor $Z_{0}$ and adaptive degradation kernel $Z_{1}$. In the first stage, DDEM processes the degraded image $I_{\text{LQ}}$ and ground-truth $I_{\text{GT}}$. Otherwise, it uses only $I_{\text{LQ}}$. The input undergoes several MOS2D Degradation-Aware Layers (MDSLs) to extract degradation feature map $F\in\mathbb{R}^{H\times W\times C_{d}}$. Subsequently, the global descriptor $Z_{0}$ and adaptive kernels $Z_{1}$ are derived from $F$ as follows:
@@ 4.2 Morton-Order 2D-Selective-Scan Module (MOS2D)
To effectively model spatially heterogeneous degradations in 2D images, our MOS2D employs Morton-Order scan, as illustrated in Fig. 1(d). This converts 2D spatial features into locality-preserving 1D sequence, facilitating structured feature interaction and aggregation by SSM.
@@
In the DDEM, Morton-order coding is followed by standard SSM operations to effectively extract global degradation descriptor $Z_{0}$ and adaptive degradation kernel $Z_{1}$. In contrast, in our main restoration backbone, the MOS2D module employs degradation-aware modulations using $Z_{0}$ and $Z_{1}$ through DAFM and DSAM, detailed in Fig. 5. This dual-modulation ensures that the MOS2D is dynamically conditioned on both long range contextually aware and spatially adaptive restoration.
Degradation-Adaptive Feature Modulation (DAFM). The $i$-th layer’s input feature map $F_{i}$ is first modulated by the global degradation descriptor $Z_{0}$. As shown in Fig. 5, $Z_{0}$ is expanded and split to produce channel-wise adaptive weights $Z_{0}^{w}$ and biases $Z_{0}^{b}$, which applied to $F_{i}$ using a feature-wise linear modulation operation, thus incorporates global degradation characteristics:
@@
4.3 Loss Function
Our model is trained in two stages. In the first stage, the loss $\mathcal{L}_{\text{stage1}}$ combines a $\mathcal{L}_{1}$ loss and a correlation loss $\mathcal{L}_{\text{cor}}$ to ensure accuracy and structural fidelity between the output of MODEM $I_{\text{HQ}}$ and ground-truth $I_{\text{GT}}$:
The correlation loss $\mathcal{L}_{\text{cor}}(I_{\text{HQ}},I_{\text{GT}})$ is based on the Pearson correlation coefficient $\rho(I_{\text{HQ}},I_{\text{GT}})$:
@@ 4.3 Loss Function
Figure 6: Visual comparison on the real world snow dataset [35]. Compared to the prior methods [78, 44, 54], where “+" denotes additional training of Histoformer, our MODEM achieves superior results without additional training, effectively removing real-world snow and restoring clean details.
@@ 5 Experimental Validation
Our MODEM is implemented in PyTorch [45] and trained on 4 NVIDIA RTX 3090 GPUs in two stages. We employ the AdamW optimizer [37] with a Cosine Annealing Restart Cyclic LR scheduler [36]. It is trained on an all-weather dataset [56, 54], consistent with prior works [30, 56, 44, 54, 78, 66]. For evaluation, MODEM is tested on established benchmarks including Test1 [29, 30], RainDrop [46], Snow100K-L/S [35], and a real-world snow test set [35]. Due to the page limit, more details are given in Sections A.1 and A.2.
@@ 5.1 Comparisons with State-of-the-arts
Quantitative Comparison. Our quantitative evaluation assesses MODEM against both unified models [30, 56, 8, 78, 44, 66, 54] and task-specific methods [58, 7, 31, 35, 73, 6, 69, 33, 20, 47, 46, 61, 55, 76, 29, 21, 68]. As reported in Table 1, MODEM achieves SOTA performance, with an average PSNR improvement of 0.5dB across benchmarks [29, 30, 46, 35], and a notable 1.02dB gain on Outdoor-Rain [29, 30]. While slightly below Histoformer on RainDrop [46] by 0.05dB PSNR, MODEM yields compelling visual results, a point further elaborated in our qualitative comparisons. Furthermore, comparisons with leading task-specific methods in Tables 3, 3 and 4 confirm MODEM consistently attains SOTA results. These evaluations underscore MODEM’s capability to effectively estimate and adapt to diverse, spatially heterogeneous weather degradations. Additionally, the comparisons of complexity can be found in Section A.3.
@@
Figure 7: T-SNE results of Histoformer [54] and MODEM, which reflect that MODEM exhibits better clustering of features corresponding to different weather types than Histoformer.

### v2

@@ 1 Introduction
Computer vision systems are increasingly integral to critical applications, such as autonomous driving [51, 1] and intelligent surveillance [49, 70], demanding reliable performance in diverse environments. However, their effectiveness deteriorates significantly under adverse weather conditions, such as rain [16, 42, 57, 75, 78, 92], haze [76, 77, 21, 62, 71, 83, 85], and snow [38, 45, 59, 87, 13, 12], which introduce complex visual artifacts and obscure critical scene information. Thus, effectively restoring clear images from such weather-degraded inputs is essential to boost the robustness and safety of modern computer vision technologies in real-world deployments.
Early task-specific methods [2, 4, 23, 60, 61, 30, 32, 90] largely rely on physical models or hand-crafted statistical priors tailored to individual weather phenomena. Due to the limited feature representation capabilities, these schemes are brittle in the face of complex scenes or deviations from assumed degradation patterns. With the advent of deep learning, numerous deep networks have achieved impressive performance by training on large-scale datasets for specific tasks such as image deraining [16, 42, 57, 75, 78, 5, 41, 69, 83], dehazing [76, 77, 21, 62, 71, 83, 85], or desnowing [38, 45, 59, 87]. These models excel at implicitly learning the inverse mapping through extensive supervision. But their highly specialized nature necessitates separate models for each weather condition, severely limiting scalability and practical deployment in unconstrained environments.
@@
Motivated by the above insight, we propose a novel Morton-Order Degradation Estimation Mechanism (MODEM) that estimates degradation state via SSM and guides adaptive restoration in degraded images. At the heart of MODEM lies the Morton-Order 2D-Selective-Scan Module (MOS2D), which integrates SSM with Morton-coded spatial traversal. Unlike conventional raster scanning, the Morton-order scan follows a hierarchical, locality-preserving sequence [50, 34, 53, 31, 7, 54] that enables structured and efficient modeling of both local and long-range dependencies, as shown in Fig. 1(d). To complement MOS2D, we further introduce a Dual Degradation Estimation Module (DDEM), designed to extract two complementary forms of degradation information from the input: (i) a global degradation representation that encapsulates high-level weather characteristics such as type and severity, and (ii) a set of spatially adaptive kernels that encode local degradation structures and variations. These dual degradation representations are then utilized to dynamically modulate the restoration process. Specifically, the global representation adaptively influences feature transformations within MODEM, while the spatially adaptive kernels guide the matrix construction within MOS2D, refining the spatial dependencies captured along the Morton-order sequence. This dual-modulation strategy equips the restoration network with both global awareness and local sensitivity, thereby delivering more precise and effective restoration. Our primary contributions can be summarized as follows:
@@ 2 Related Work
Raindrop Removal. Early efforts addressed specific scenarios [15] or adherent raindrops in videos [80]. Deep learning subsequently provided more generalizable solutions, including learning from synthetic photorealistic data [22], using attention-guided GANs for realistic inpainting [57], and dedicated networks for visibility through raindrop-covered windows [58]. General restoration architectures like Dual Residual Networks [43] and Adaptive Sparse Transformers [92] are also applicable, with advanced dual attention-in-attention models [86] tackling joint rain streak and raindrop removal.
@@
Snow Removal. While specific traditional priors for snow removal are less common compared to other weather conditions, principles from general image restoration were often adapted. Deep learning solutions have gained traction, including powerful general-purpose backbones [13, 12] and more specialized networks. Desnowing-specific models often involve context-aware designs [45] or leverage semantic and depth priors [87]. Techniques inspired by classical matrix decomposition [59] and methods for online video processing [38] have also been integrated into deep frameworks.
@@
Figure 3: With respect to a sample (a), (b)-(d) visualize the long-range $CAh_{k-1}$, (c) local $CBx_{k}$, and (d) output of MOS2D $y_{k}$, respectively. More cases can be found in Appendix A.4
@@ 3 Bridging Image Degradation Estimation and State Space Modeling
By comparing Eq.˜1 with Eq.˜3, we interpret SSM components as degradation estimators, as illustrated in Fig.˜2. The term $CAh_{k-1}$ conceptually models the long-range, spatially propagated degradation context, as shown in Fig.˜3(b). The previous hidden state $h_{k-1}$ summarizes accumulated degradation cues (e.g., overall haze level, rain intensity) along the Morton-order scan path. The state transition matrix $A$ then models how these summarized characteristics evolve or persist as the scan progresses spatially (e.g., the smooth attenuation of visibility due to widespread haze or the consistent statistical properties of rain streaks over a larger area.). It essentially dictates “how the long-range state evolves over pixels” as noted in Fig.˜2. The output matrix $C$ then translates this evolved context $Ah_{k-1}$ into a contribution to the current output $y_{k}$, reflecting broader, non-local degradation patterns.
Concurrently, the $CBx_{k}$ term accounts for the impact of the immediate, local input features $x_{k}$, as visualized in Fig.˜3(c). The input matrix $B$ determines “how the local information affects the current state” (Fig.˜2), allowing features from $x_{k}$ (e.g., local degradation artifacts like dense fog patches, or crucial local image content like fine textures/sharp edges) to directly influence the current hidden state $h_{k}$. This enables the SSM to react to fine-grained local evidence, with $C$ projecting this locally-informed component into $y_{k}$ for specific adjustments based on immediate observations.
Figure 4: (a) Overall architecture of MODEM. (b) The DDEM for extracting global descriptor $Z_{0}$ and adaptive degradation kernel $Z_{1}$ degradation priors. (c) The MDSL incorporating the core MOS2D module (d) within a residual block. The blue-colored components indicate elements exclusive to the first training stage. NN, $M_{1}$, $M_{2}$ denote the number of the corresponding module, respectively.
@@ 4 Morton-Order Degradation Estimation Mechanism
The Morton-Order Degradation Estimation Mechanism (MODEM), depicted in Fig.˜4(a), employs a two-stage training strategy designed to accurately learn degradation characteristics.
Stage 1: In the first stage, the Dual Degradation Estimation Module (DDEM), shown in Fig.˜4(b), is provided with both the degraded image $I_{\text{LQ}}$ and its corresponding ground-truth $I_{\text{GT}}$ from the training set. This allows the DDEM to explicitly learn the mapping from an image pair to its underlying degradation patterns. It outputs two degradation priors: a global descriptor $Z_{0}$ and a spatially adaptive degradation kernel $Z_{1}$. These priors, representing the degradation information, are then injected into MODEM’s main restoration backbone. The backbone itself, comprised of $N$ stacked MOS2D Degradation-Aware Layers (MDSLs) detailed in Fig.˜4(c), only receives the degraded image $I_{\text{LQ}}$ and uses these priors to perform adaptive restoration. Each MDSL leverages the Morton-Order 2D-Selective-Scan module (MOS2D), shown in Fig.˜4(d) and Fig.˜5, to adaptively modulate features based on the degradation priors $Z_{0}$ and $Z_{1}$. The degradation representation learned by the DDEM in this stage serves as the supervisory target for the Stage 2.
Stage 2: In the second stage, the inputs to both the DDEM and the main backbone consist of only the degraded image $I_{\text{LQ}}$, without any GT. The Stage 2 model inherits its parameters from Stage 1, and the DDEM in this stage receives only the degraded image $I_{\text{LQ}}$ as input. Simultaneously, there is a frozen DDEM receiving both the GT and the degraded image. The degradation information from this frozen DDEM is then used to supervise the trainable DDEM.
@@ 4.1 Dual Degradation Estimation Module (DDEM)
The Dual Degradation Estimation Module (DDEM), shown in Fig.˜4(b), extracts global degradation descriptor $Z_{0}$ and adaptive degradation kernel $Z_{1}$. In the first stage, DDEM processes the degraded image $I_{\text{LQ}}$ and ground-truth $I_{\text{GT}}$. Otherwise, it uses only $I_{\text{LQ}}$. The input undergoes several MOS2D Degradation-Aware Layers (MDSLs) to extract degradation feature map $F\in\mathbb{R}^{H\times W\times C_{d}}$. Subsequently, the global descriptor $Z_{0}$ and adaptive kernels $Z_{1}$ are derived from $F$ as follows:
@@ 4.2 Morton-Order 2D-Selective-Scan Module (MOS2D)
To effectively model spatially heterogeneous degradations in 2D images, our MOS2D employs Morton-Order scan, as illustrated in Fig.˜1(d). This converts 2D spatial features into locality-preserving 1D sequence, facilitating structured feature interaction and aggregation by SSM.
@@
In the DDEM, Morton-order coding is followed by standard SSM operations to effectively extract global degradation descriptor $Z_{0}$ and adaptive degradation kernel $Z_{1}$. In contrast, in our main restoration backbone, the MOS2D module employs degradation-aware modulations using $Z_{0}$ and $Z_{1}$ through DAFM and DSAM, detailed in Fig.˜5. This dual-modulation ensures that the MOS2D is dynamically conditioned on both long range contextually aware and spatially adaptive restoration.
Degradation-Adaptive Feature Modulation (DAFM). The $i$-th layer’s input feature map $F_{i}$ is first modulated by the global degradation descriptor $Z_{0}$. As shown in Fig.˜5, $Z_{0}$ is expanded and split to produce channel-wise adaptive weights $Z_{0}^{w}$ and biases $Z_{0}^{b}$, which applied to $F_{i}$ using a feature-wise linear modulation operation, thus incorporates global degradation characteristics:
@@
Figure 6: Visual comparison on the real world snow dataset [45]. Compared to the prior methods [95, 55, 65], where $\dagger$}" denotes the Histoformer [65] for real snow from their official repository, our MODEM achieves superior results without additional training.
@@
4.3 Loss Function
Our model is trained in two stages. In the first stage, the loss $\mathcal{L}_{\text{stage1}}$ combines a $\mathcal{L}_{1}$ loss and a correlation loss $\mathcal{L}_{\text{cor}}$ [65] to ensure accuracy and structural fidelity between the output of MODEM $I_{\text{HQ}}$ and ground-truth $I_{\text{GT}}$:
The correlation loss $\mathcal{L}_{\text{cor}}(I_{\text{HQ}},I_{\text{GT}})$ [65] is based on the Pearson correlation coefficient $\rho(I_{\text{HQ}},I_{\text{GT}})$:
@@ 5 Experimental Validation
Our MODEM is implemented in PyTorch [56] and trained on 4 NVIDIA RTX 3090 GPUs in two stages. We employ the AdamW optimizer [47] with a Cosine Annealing Restart Cyclic LR scheduler [46]. It is trained on an all-weather dataset [67, 65], consistent with prior works [40, 67, 55, 65, 95, 79]. For evaluation, MODEM is tested on established benchmarks including Test1 [39, 40], RainDrop [57], Snow100K-L/S [45], and a real-world snow test set [45]. Due to the page limit, more details are given in Appendix A.1 and A.2.
@@ 5.1 Comparisons with State-of-the-arts
Quantitative Comparison. Our quantitative evaluation assesses MODEM against both unified models [40, 67, 11, 95, 55, 79, 65] and task-specific methods [69, 10, 41, 45, 87, 9, 82, 43, 27, 58, 57, 73, 66, 93, 39, 29, 81]. As reported in Table˜1, MODEM achieves SOTA performance, with an average PSNR improvement of 0.5dB across benchmarks [39, 40, 57, 45], and a notable 1.02dB gain on Outdoor-Rain [39, 40]. While slightly below Histoformer on RainDrop [57] by 0.05dB PSNR, MODEM yields compelling visual results, a point further elaborated in our qualitative comparisons. Furthermore, comparisons with leading task-specific methods in Tables˜3, 3 and 4 confirm MODEM consistently attains SOTA results. These evaluations underscore MODEM’s capability to effectively estimate and adapt to diverse, spatially heterogeneous weather degradations. Additionally, the comparisons of complexity can be found in Appendix A.3.
@@
Figure 7: T-SNE results of Histoformer [65] and MODEM, which reflect that MODEM exhibits better clustering of features corresponding to different weather types than Histoformer [65].
Qualitative comparison. We provide qualitative comparisons across diverse scenarios in Figs.˜8 and 6. For image desnowing on Snow100K [45], MODEM effectively removes heavy snowflakes and visual artifacts that other models [95, 55, 65] struggle to address. For joint deraining and dehazing on Outdoor-Rain [39, 40], MODEM excels in restoring richer texture details and produces images with noticeably fewer artifacts. For raindrop removal on the Raindrop [57], MODEM again demonstrates superior detail preservation and artifact reduction. Furthermore, on real-world snowy images [45], MODEM achieves superior results as shown in Fig.˜6 even without any additional fine-tuning, showcasing excellent generalization and real-world applicability, which can be attributed to its profound understanding and adaptive modeling of degradation characteristics. Due to the page limit, more visual comparisons can be found in Appendix A.8.
5.2 Perceptual Quality and Real-World Performance
Comparison of perceptual metrics. We report referenced (LPIPS [88]) and non-referenced (Q-Align [72], MUSIQ [33]) scores, all computed using the pyiqa library [6], with the same settings applied for all methods. As shown in Table˜5, our method achieves SOTA perceptual quality.
Comparison of real-world datasets. To provide quantitative evidence of our model’s performance on the challenging real-world scenarios, we evaluated our model on the real-world data from the Snow100K-Real [45], RainDrop [57], RESIDE [36], NTURain [8] and WeatherStream [84] datasets using the Q-Align [72]. The results are presented in Table˜6. As the results show, our method achieves state-of-the-art performance on real-world datasets when compared to the previous state-of-the-art method, Histoformer [65], and the diffusion-based approach, WeatherDiff [55].
[…]
Method Snow100K-L Snow100K-S Outdoor Raindrop Snow100K-Real
LPIPS
Histoformer [65] 0.0919 0.0445 0.0778 0.0672
WeatherDiff [55] 0.0982 0.0541 0.0887 0.0615
MODEM (Ours) 0.0880 0.0407 0.0699 0.0650
Q-Align
Histoformer [65] 3.7207 3.7598 4.1445 4.0156 3.5449
WeatherDiff [55] 3.4531 3.5293 3.8691 4.0000 3.4512
MODEM (Ours) 3.7324 3.7695 4.1875 4.0664 3.5586
MUSIQ
Histoformer [65] 64.2526 64.2581 67.7461 68.4852 59.4040
WeatherDiff [55] 62.6267 63.1278 67.4814 69.3608 59.4493
MODEM (Ours) 64.2438 64.2853 68.2926 69.7925 59.6042
Table 6: Comparison of different methods on various real-world datasets using the Q-Align metric.
Method Snow100K-Real RainDrop NTURain RESIDE WeatherStream
WeatherDiff [55] 3.4531 4.0000 3.2031 3.4219 1.9561
Histoformer [65] 3.7207 4.0156 3.2266 3.2891 1.9434
MODEM (Ours) 3.7324 4.0664 3.2891 3.3164 1.9863
5.3 Ablation Studies
We evaluate the impact of the Morton-Order scan, DDEM, DAFM and DSAM. Morton-Order scan facilitates structured spatial feature processing within the SSM, enabling the model to better capture contextual dependencies and preserve local structural coherence. As shown in Table˜7, configurations incorporating the Morton scan generally yield improved performance, underscoring its benefit in organizing spatial information for sequential modeling. DDEM provides the degradation priors $Z_{0}$ and $Z_{1}$ that guide the DAFM and DSAM. As indicated in Table˜7, configurations lacking DDEM exhibit a significant performance drop. DAFM, which utilizes the global degradation prior $Z_{0}$, plays a crucial role in adaptively modulating features based on the overall estimated weather type and severity. DSAM, guided by the spatially adaptive kernel $Z_{1}$, allows the MOS2D to selectively focus on and modulate features pertinent to local degradation characteristics or specific image regions. While removing DSAM shows marginal gains on certain snow metrics, it impairs performance on tasks like raindrop removal, which demands strong local adaptation due to the highly local nature of raindrops, and on mixed rain&haze scenarios where intricate local texture recovery is challenging. Thus, DSAM, guided by the spatially adaptive kernel $Z_{1}$, is crucial for achieving robust, balanced performance across diverse weather conditions by guiding the network to selectively address these local characteristics. The combination of all components employed in full MODEM, consistently yields the best results. This result is further supported by t-SNE [68] in Fig.˜7. Compared to Histoformer [65], MODEM exhibits significantly improved clustering of features corresponding to different weather types. This indicates that the combined action of the components enables MODEM to learn more discriminative and well-separated feature representations, which directly contributes to enhance restoration capabilities. More ablation studies can be found in Appendix A.6.